OpenStudy (anonymous):

Which of the following is a zero for the function f(x) = (x – 15)(x + 1)(x – 10)? x = –15 x = –10 x = 1 x = 15

OpenStudy (anonymous):

Can you help me?? @Meepi

OpenStudy (amistre64):

what do we know about multiplication, that gets us a zero?

OpenStudy (amistre64):

6, times what ... equals zero? 3, times what .... equals zero? 198726735544783836734, times what .... equals zero?

OpenStudy (anonymous):

6 times 0 = zero

OpenStudy (anonymous):

I dont get what youre trying to say by that @amistre64

OpenStudy (amistre64):

its a property, that is useful to us here. you have 3 numbers, multiplied together. (x – 15), (x + 1), and (x – 10) to determine what value of x, creates a zero.... we only need to define what values of x will produce a zero in any of the numbers given

OpenStudy (amistre64):

(x – 15) = 0, what is x? (x + 1) = 0, what is x? (x – 10) = 0, what is x?

OpenStudy (anonymous):

x always zero??? @amistre64

OpenStudy (amistre64):

lets see, if x = 0 then we get: (0 – 15) = -15, not it (0 + 1) = 1 , not it (0 – 10) = -10 , not it when x=0, we do not produce a number that is zero.

OpenStudy (anonymous):

but it's asking for the function f(x) = (x – 15)(x + 1)(x – 10)? Sorry im confused... @amistre64

OpenStudy (amistre64):

do you see that when "x" is equal to a certain value, that we get 3 numbers that multiply together?

OpenStudy (anonymous):

So its not zero, would it be x= 1 since the number would than equal its self

OpenStudy (amistre64):

f(x) = (x – 15)(x + 1)(x – 10) if x = 0 then we get: f(0) = (0 – 15)(0 + 1)(0 – 10) f(0) = (–15)(1)(–10) = 150 since 150 is not equal to zero, then x=0 is does not create a zero.

OpenStudy (amistre64):

lets try x=1 f(1) = (1 – 15)(1 + 1)(1 – 10) not0 not0 not0 f(1) = (–14)(2)(–9) = not0 recall the property i hinted at. to multiply numbers together and get zero, one or more of the numbers HAVE to be zero.

OpenStudy (amistre64):

f(x) = (x – 15)(x + 1)(x – 10) we can take each number they give us, and equal it to zero (x – 15) = 0, what would x have to be for this to be zero? (x + 1) = 0, what would x have to be for this to be zero? (x – 10) = 0, what would x have to be for this to be zero?

OpenStudy (anonymous):

15 -1 10 I'm still confused lol

OpenStudy (amistre64):

you are absolutely correct tho :)

OpenStudy (amistre64):

f(15) = (15 – 15)(15 + 1)(15 – 10) f(15) = 0 (15 + 1) (15 – 10) = 0 and it works for the other zeros you found as well

OpenStudy (anonymous):

So what would the answer be?

OpenStudy (amistre64):

so given our set of zeros: x = {15,-1,10} which of the options will be correct? x = –15 x = –10 x = 1 x = 15

OpenStudy (anonymous):

15???

OpenStudy (amistre64):

correct :)

OpenStudy (anonymous):

Thank you!